# Latent Diffusion Models with Masked AutoEncoders

### v1

@@ Abstract
In spite of remarkable potential of the Latent Diffusion Models (LDMs) in image generation, the desired properties and optimal design of the autoencoders have been underexplored. In this work, we analyze the role of autoencoders in LDMs and identify three key properties: latent smoothness, perceptual compression quality, and reconstruction quality. We demonstrate that existing autoencoders fail to simultaneously satisfy all three properties, and propose Variational Masked AutoEncoders (VMAEs), taking advantage of the hierarchical features maintained by Masked AutoEncoder. We integrate VMAEs into the LDM framework, introducing Latent Diffusion Models with Masked AutoEncoders (LDMAEs). Through comprehensive experiments, we demonstrate significantly enhanced image generation quality and computational efficiency.
@@ 1 Introduction
Despite the notable success of the additional loss terms in autoencoders, their specific roles in shaping the characteristics of autoencoders, as well as their impact on the entire LDM framework, remain underexplored. A fundamental question arises: what properties should an autoencoder possess to effectively support LDMs? Since the choice of autoencoder directly influences the quality and efficiency of LDMs, this is a crucial question for designing an autoencoder. To address this question, we propose three key properties that an autoencoder should satisfy to effectively compress the data, preserving essential semantics, and eventually to facilitate high-quality image generation in the LDM framework: 1) latent smoothness, 2) perceptual compression quality, and 3) reconstruction quality.
First, latent smoothness ensures that minor perturbations in the latent representation does not drastically change the generated output. A sufficiently smooth latent space allows the model to generate high-quality images robustly even when the diffusion model predicts the exact latent code with some error. We reveal through an empirical analysis on the stability of autoencoder embeddings under small perturbations (Sec. 5.1) that autoencoders with deterministic encoding, such as vanilla AutoEncoders (AEs), lack this property, whereas probabilistic encoding, as used in VAEs, much better satisfies this requirement.
Second, perceptual compression quality measures how effectively an autoencoder encodes the images into a simplified latent space, maximizing the reduction of perceptual details while preserving semantics as much as possible. However, the boundary between ‘semantics’ and ‘perceptual details’ is inherently ambiguous, as they reside on a continuous spectrum. This spectrum ranges from pixel-level (minimal compression) to object-level (maximum compression). In Sec. 5.1, we observe that SD-VAEs aggressively compress input images, resulting in compression close to the object-level. This simplifies the latent space, making it easier for diffusion models to predict latent codes, but sacrifices fine-grained details. To mitigate this, we propose hierarchical compression, which organizes the latent space into a continuous hierarchy: first grouping images into object-level clusters, then subdividing each object into part-level clusters (e.g., facial features), and finally refining them into fine-detail clusters. This structure preserves global simplicity while retaining local details, effectively balancing compression and reconstruction quality.
@@
From systematical assessments of these properties (see Sec. 5.1 for details), we provide a concise summary of various autoencoders in Fig. 2. We clearly observe that no existing autoencoders satisfy all of these desired properties, lacking at least one of them. Inspired by the recent findings of masked autoencoding [34] that it construct a well-aligned hierarchical feature space, we propose a novel transformer-based Variational Masked AutoEncoders (VMAEs). VMAE achieves both latent smoothness and high reconstruction quality through probabilistic alignment, tailored loss design, and structural enhancements, while retaining the hierarchical compression benefits of MAEs. Additionally, VMAE significantly improves computational efficiency, using only 13.4% of the parameters and 4.1% of the GFLOPs of SD-VAE. It also converges 2.7 times faster with greater training stability.
@@ 4.1 Support of Latent Space
We emphasize that probabilistic latent space is crucial for training LDMs, as it ensures a smooth latent space—a fundamental requirement of the diffusion process. An overly sparse or low-dimensional latent manifold can hinder diffusion modeling by causing inconsistent score matching [38].
@@ 4.2 Perceptual Compression
Fig. 4 (top) illustrates unsupervised clustering of the local features of SD-VAE for a giraffe image, with various number of clusters. While SD-VAE clearly distinguishes the main object (giraffe) from the background (2 clusters), it struggles to capture finer details like fur patterns (4 clusters). This suggests that its compression ability is limited to the highest object-level, renouncing to represent finer details in the latent space. This simplified latent space can facilitate the learning process of the diffusion model [2], but it may result in the loss of perceptual details at decoding, as most fine details are discarded during the encoding process. Also, as the fine-grained features are entangled in the latent space, downstream tasks like image editing, where disentangled features are particularly critical [25, 12, 35] would not be performed well. Based on these intuitions, we explore an alternative autoencoder with optimal perceptual compression.
@@ 5 Evaluation on Autoencoders
In this section, we first evaluate our VMAEs in the aspect of latent space, comparing with other autoencoders introduced in Sec. 3.1, i.e., vanilla AutoEncoders (AEs) [31], Denoising AutoEncoders (DAEs) [42], Variational AutoEncoders (VAEs) [18] and StableDiffusion3 VAEs (SD-VAEs) [10]. All autoencoders are trained on the ImageNet-1K [6] training set with their respective optimal hyperparameters using 8 NVIDIA A100 GPUs (40GB). Elaboration on datasets and implementation details can be found in Sec. B.1.
@@ 5.1 Results and Analysis
Latent Space Smoothness. As introduced in Sec. 4.1, the latent space should be smooth and continuous rather than consisting of sparse points to ensure stable and consistent score matching. In the LDMs framework, this property is closely related to the difficulty of training the diffusion model. As illustrated in Fig. 3, deterministic encoders (AE, DAE) project inputs into discrete and sparse latent points, restricting the decodable space to the discrete isolated points. In contrast, probabilistic encoders (VAE, SD-VAE, VMAE) map inputs to latent distributions, forming a continuous latent space that broadens the decodable region. A continuous space is preferable to discrete points for diffusion models to learn, as it allows generations within the vicinity of target points rather than requiring exact matches.
To assess the smoothness of the latent space, we compare the robustness of each autoencoder to a noise with varied perturbation ratios, assuming that a broader decodable (supported) space would be more robust to latent perturbations and thus lead to a better quality of decoded images. Fig. 6 reports rFID scores of each autoencoder with varied noise levels, measuring the quality of generated images. First, the deterministic autoencoders (AE and DAE) show reasonably good performance when the noise level is extremely low, while the rFID score significantly drops when a larger level of perturbation is applied. This suggests that deterministic encoding approaches would destabilize the diffusion model’s learning process due to their limited support in the latent space. In spite of their completely deterministic design, they still show robustness for a very small noise, probably because the training process causes some level of errors when they map images to the latent space and reconstruct them, and the model learns to robustly map within this small range of noise. Second, VAE shows minimal increase in rFID across perturbation ratios, indicating that its decodable space is most widely supported. Nevertheless, its high absolute rFID scores make it less suitable for LDMs. Lastly, SD-VAE and VMAE demonstrate combined strengths, achieving the best quality (in rFID) while more robust to noise compared to AE and DAE. This indicates that a probabilistic latent space indeed results in a larger decodable space which is more desirable for LDMs. Comparing the two, our VMAE consistently outperforms the strongest baseline, SD-VAE across all noise levels. This conclusion is consistent with the actual generation quality, demonstrated in Tab. 3.
@@
Model Efficiency. We compare the model cost in Tab. 2. Our VMAE demonstrates significantly faster training and inference, with considerably a smaller model size. Faster inference is particularly favorable for training LDMs, as image encoding is required at each iteration.
@@ 6.1 Results and Analysis
AE and DAE, which lack smooth latent spaces, show weaker generation performance than probabilistic models like SD-VAE and VMAE, which benefit from smoother latent representations. Among deterministic models, DAE performs slightly better than AE, indicating that DAE’s latent space is actually more robust [42].
VAE achieves the smoothest latent space, but its poor reconstruction quality results in the worst generation scores on both datasets. This highlights that latent smoothness alone is insufficient without reliable reconstruction. Notably, the gap between VAE and deterministic models (AE, DAE) is much smaller on CelebA-HQ than on ImageNet-1K. This is likely because the ImageNet-1K dataset is often poorly preprocessed, leaving objects off-center or smaller, with large, noisy backgrounds. Such background noise reduces sensitivity to fine reconstruction errors, especially in perceptual metrics like FID. In contrast, CelebA-HQ images are tightly cropped, with large, centered faces and minimal background, making them more sensitive to small errors in the predicted latent codes, which can be resolved by a sufficiently smooth latent space.
SD-VAE balances all three properties well, achieving solid performance on both datasets. VMAE further improves each property and consistently achieves the best scores across all metrics. Moreover, VMAE’s superior pixel-level reconstruction quality, while not directly reflected in FID and IS, likely enhances the perceptual quality of generated images. Figs. 1 and 9 shows generated samples from ImageNet-1K and CelebA-HQ, demonstrating high-resolution results with balanced quality across diverse objects and faces.
@@ B.1 Experimental Setup for Autoencoders
Implementation Details. We train all autoencoders using the optimal hyperparameters for each model on 8 NVIDIA A100 GPUs (40GB). The base learning rate is set to $10^{-5}$ for convolution-based models (AEs, DAEs, VAEs, and SD-VAEs), while $10^{-4}$ for our VMAEs. Global batch size is set to 2048 except for the SD-VAEs, which require smaller batch size (256) for stable training of the adversarial network, following the implementation in VQGAN [9].

### v2

@@ Abstract
In spite of the remarkable potential of Latent Diffusion Models (LDMs) in image generation, the desired properties and optimal design of the autoencoders have been underexplored. In this work, we analyze the role of autoencoders in LDMs and identify three key properties: latent smoothness, perceptual compression quality, and reconstruction quality. We demonstrate that existing autoencoders fail to simultaneously satisfy all three properties, and propose Variational Masked AutoEncoders (VMAEs), taking advantage of the hierarchical features maintained by Masked AutoEncoders. We integrate VMAEs into the LDM framework, introducing Latent Diffusion Models with Masked AutoEncoders (LDMAEs). Through comprehensive experiments, we demonstrate significantly enhanced image generation quality and computational efficiency.
@@ 1 Introduction
Despite the notable success of the additional loss terms in autoencoders, their specific roles in shaping the characteristics of autoencoders, as well as their impact on the entire LDM framework, remain underexplored. A fundamental question arises: what properties should an autoencoder possess to effectively support LDMs? Since the choice of autoencoder directly influences the quality and efficiency of LDMs, this is a crucial question for designing an autoencoder. To address this question, we propose three key properties that an autoencoder should satisfy to effectively compress data, preserve essential semantics, and eventually facilitate high-quality image generation in the LDM framework: 1) latent smoothness, 2) perceptual compression quality, and 3) reconstruction quality.
First, latent smoothness ensures that minor perturbations in the latent representation do not drastically change the generated output. A sufficiently smooth latent space allows the model to generate high-quality images robustly, even when the diffusion model predicts latent codes with some error. We reveal through an empirical analysis on the stability of autoencoder embeddings under small perturbations (Sec. 5.1) that autoencoders with deterministic encoding, such as vanilla AutoEncoders (AEs), lack this property, whereas probabilistic encoding, as used in VAEs, much better satisfies this requirement.
Second, perceptual compression quality measures how effectively an autoencoder encodes the images into a simplified latent space, maximizing the reduction of perceptual details while preserving semantics as much as possible. However, the boundary between ‘semantics’ and ‘perceptual details’ is inherently ambiguous, as they exist on a continuous spectrum. This spectrum ranges from pixel-level (minimal compression) to object-level (maximum compression). In Sec. 5.1, we observe that SD-VAEs aggressively compress input images, resulting in compression close to the object-level. This simplifies the latent space, making it easier for diffusion models to predict latent codes, but sacrifices fine-grained details. To mitigate this, we propose hierarchical compression, which organizes the latent space into a continuous hierarchy: first grouping images into object-level clusters, then subdividing each object into part-level clusters (e.g., facial features), and finally refining them into fine-detail clusters. This structure preserves global simplicity while retaining local details, effectively balancing compression and reconstruction quality.
@@
From systematical assessments of these properties (see Sec. 5.1 for details), we provide a concise summary of various autoencoders in Fig. 2. We clearly observe that no existing autoencoders satisfy all of these desired properties, lacking at least one of them. Inspired by the recent findings of masked autoencoding [36] that it constructs a well-aligned hierarchical feature space, we propose a novel transformer-based Variational Masked AutoEncoders (VMAEs). VMAE achieves both latent smoothness and high reconstruction quality through probabilistic alignment, tailored loss design, and structural enhancements, while retaining the hierarchical compression benefits of MAEs. Additionally, VMAE significantly improves computational efficiency, using only 13.4% of the parameters and 4.1% of the GFLOPs of SD-VAE. It also converges 2.7 times faster with greater training stability.
@@ 4.1 Support of Latent Space
We emphasize that probabilistic latent space is crucial for training LDMs, as it ensures a smooth latent space—a fundamental requirement of the diffusion process. An overly sparse or low-dimensional latent manifold can hinder diffusion modeling by causing inconsistent score matching [4, 40].
@@ 4.2 Perceptual Compression
Fig. 4 (top) illustrates unsupervised clustering of the local features of SD-VAE for a giraffe image, with various number of clusters. While SD-VAE clearly distinguishes the main object (giraffe) from the background (2 clusters), it struggles to capture finer details like fur patterns (4 clusters). This suggests that its compression ability is limited to the highest object-level, renouncing to represent finer details in the latent space. This simplified latent space can facilitate the learning process of the diffusion model [2], but it may result in the loss of perceptual details at decoding, as most fine details are discarded during the encoding process. Also, as the fine-grained features are entangled in the latent space, downstream tasks like image editing, where disentangled features are particularly critical [27, 14, 37, 12] would not be performed well. Based on these intuitions, we explore an alternative autoencoder with optimal perceptual compression.
@@ 5 Evaluation on Autoencoders
In this section, we first evaluate our VMAEs in the aspect of latent space, comparing with other autoencoders introduced in Sec. 3.1. All autoencoders are trained on the ImageNet-1K [7] training set with their respective optimal hyperparameters using 8 NVIDIA A100 GPUs (40GB). Elaboration on datasets and implementation details can be found in Sec. B.1.
@@ 5.1 Results and Analysis
Latent Space Smoothness. As introduced in Sec. 4.1, the latent space should be smooth and continuous rather than consisting of sparse points to ensure stable and consistent score matching. As illustrated in Fig. 3, deterministic encoders (AE, DAE) project inputs into discrete and sparse latent points, restricting the decodable space to the discrete isolated points. In contrast, probabilistic encoders (VAE, SD-VAE, VMAE) map inputs to latent distributions, forming a continuous latent space that broadens the decodable region. A continuous space is preferable to discrete points for diffusion models to learn, as it allows generations within the vicinity of target points rather than requiring exact matches.
To assess the smoothness of the latent space, we compare the robustness of each autoencoder to a noise with varied perturbation ratios. Fig. 6 reports rFID scores of each autoencoder with varied noise levels, measuring the quality of generated images. First, the deterministic autoencoders (AE and DAE) show reasonably good performance when the noise level is extremely low, while the rFID score significantly drops when a larger level of perturbation is applied. This suggests that deterministic encoding approaches would destabilize the diffusion model’s learning process due to their limited support in the latent space. In spite of their completely deterministic design, they still show robustness for a very small noise, probably because the training process causes some level of errors when they map images to the latent space and reconstruct them, and the model learns to robustly map within this small range of noise. Second, VAE shows minimal increase in rFID across perturbation ratios, indicating that its decodable space is most widely supported. Nevertheless, its high absolute rFID scores make it less suitable for LDMs. Lastly, SD-VAE and VMAE demonstrate combined strengths, achieving the best quality (in rFID) while more robust to noise compared to AE and DAE. This indicates that a probabilistic latent space indeed results in a larger decodable space which is more desirable for LDMs. Comparing the two, our VMAE consistently outperforms the strongest baseline, SD-VAE across all noise levels. This conclusion is consistent with the actual generation quality, demonstrated in Tab. 3.
@@
Model Efficiency. We compare the model cost in Tab. 2. Our VMAE demonstrates significantly faster training and inference, with considerably smaller model size. Faster inference is particularly favorable for training LDMs, as image encoding is required at each iteration.
@@ 6.1 Results and Analysis
AE and DAE, which lack smooth latent spaces, show weaker generation performance than probabilistic models like SD-VAE and VMAE, which benefit from smoother latent representations. Among deterministic models, DAE performs slightly better than AE, indicating that DAE’s latent space is actually more robust [44]. VAE achieves the smoothest latent space, but its poor reconstruction quality results in the worst generation scores on both datasets. This highlights that latent smoothness alone is insufficient without reliable reconstruction. Notably, the gap between VAE and deterministic models (AE, DAE) is much smaller on CelebA-HQ than on ImageNet-1K. This is likely because the ImageNet-1K dataset is often poorly preprocessed, leaving objects off-center or smaller, with large, noisy backgrounds. Such background noise reduces sensitivity to fine reconstruction errors, especially in perceptual metrics like FID. In contrast, CelebA-HQ images are tightly cropped, with large, centered faces and minimal background, making them more sensitive to small errors in the predicted latent codes, which can be resolved by a sufficiently smooth latent space. SD-VAE balances all three properties well, achieving solid performance on both datasets. VMAE further improves each property and consistently achieves the best scores across all metrics. Moreover, VMAE’s superior pixel-level reconstruction quality, while not directly reflected in FID and IS, likely enhances the perceptual quality of generated images. Figs. 1 and 9 show generated samples from ImageNet-1K and CelebA-HQ, demonstrating high-resolution results with balanced quality across diverse objects and faces.
@@ B.1 Experimental Setup for Autoencoders
Implementation Details. Our VMAE adopts a symmetric ViT‐based encoder–decoder architecture. The encoder partitions an input image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ into patch tokens ${\mathbf{x}_{i}\in\mathbb{R}^{h\times w\times d}}$ with patch size $(h,w)=(32,32)$ and embedding dimension $d=192$. A masking ratio of 0.6 is applied, and the visible patches are processed by 12 Transformer layers. The decoder prepends learnable mask tokens to the encoded sequence and reconstructs the full image using another 12 Transformer layers. We impose a KL‐divergence loss on the latent representations and employ both pixel‐wise reconstruction loss and perceptual loss on the decoder outputs.
We train all autoencoders using the optimal hyperparameters for each model on 8 NVIDIA A100 GPUs (40GB). The base learning rate is set to $10^{-5}$ for convolution-based models (AEs, DAEs, VAEs, and SD-VAEs), while $10^{-4}$ for our VMAEs. Global batch size is set to 2048 except for the SD-VAEs, which require smaller batch size (256) for stable training of the adversarial network, following the implementation in VQGAN [10].